OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (1, 3) m = 2

OpenStudy (studygurl14):

y = mx + b m = 2 y = (2)x + b (1, 3) (x, y) (3) = 2(1) + b Solve for b, then plug into y= 2x + b